Word problems drive me crazy, can you show me how to solve this one?
Soybean meal is 16% protein; cornmeal is 8% protein. How many pounds of each should be used to get a 320 pound mixture that is 12% protein?

Answer by Alan3354(69443)

12% is the average, halfway between 8% and 16%, so it's 160 lb of each.
--------------------
To do it the hard way:
S + C = 320 (total pounds)
0.16S + 0.08C = 0.12*320 (total protein)
16S + 8C = 12*320
2S + C = 480 no fractions and reduced
1S + C = 320
------------- Subtract
S = 160
c = 160